twice the sum of a number and 8 is equal to three times the difference of the number and three. find the number
Answer by mala_kumarkos@yahoo.com(10)

let the number be X
therefore 2(X+8)=3(X-3)
2x+16=3x-9
16+9=3x-2x
25=x